OpenStudy (anonymous):

Write an equation for the line in slope-intercept form, given that the line passes through the point (2,3) and (3,4). Answer y = -x + 1 y = x + 1 y = x – 1 y = 2x + 1

OpenStudy (anonymous):

@privetek

OpenStudy (anonymous):

I gir B.

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) =(4-3)/(3-2) = 1 y-y1=m(x-x1) -- pick one of the points y-3=1(x-2) y=x-2+3 y=x+1